Solve in Excel

You are offered an investment with the following conditions:

  • The cost of the investment is 1,000
  • The investment pays out a sum X at the end of the first year; this payout grows at the rate of 10% per year for 11 years.

If your discount rate is 15%, calculate the smallest X which would entice you to purchase the asset.
Answer #1

Cost = PV of the payouts  
1000 = X/(1+15%) + X*(1+10%)/(1+15%)2 .......+ X*(1+10%)11/(1+15%)12
1000= X/1.15 (1 + 1.1/1.5 + (1.1/1.5)2 ....+ (1.1/1.15)10
1000 = X/1.15 *(1-(1.1/1.15)11/(1-1.1/1.5)
X = 1150/(1-(1.1/1.15)11/(1-1.1/1.5) = 1150/7.734838 = 148.68

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well